# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ Appendix B Proofs for the actor
The proof for Proposition 3.1 is the same as [60, Proposition 1], where we show
@@
We decompose the derivative of $\mathcal{L}_{a}^{\tau}=J^{\mu^{\tau}}[\alpha^{\tau}]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]$ into two terms
@@
addressing the dependence on $\alpha^{\tau}$ and $\mu^{\tau}$ respectively.
@@
$=-\beta_{a}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left\langle\nabla_{\alpha}H\left(t,x,\mu^{\tau}_{t},\alpha^{\tau}(t,x),-\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x)\right),\right$.
$\hskip 72.26999pt\nabla_{\alpha}H\left(t,x,\mu^{\tau}_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x)\right)\Big\rangle\,\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\,\mathrm{d}x\,\mathrm{d}t$
@@
We will show $(a\text{III})\geq c\mathcal{L}_{a}^{\tau}$ first. We start with a technical definition. For any $\tau\geq 0$, we define the local optimal control $\alpha^{\tau,\diamond}$ as
$\alpha^{\tau,\diamond}(t,x):=\operatorname*{arg\,max}H(t,x,\mu_{t}^{\tau},\cdot,-\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x),-\nabla_{x}^{2}V^{\mu^{\tau},\alpha^{\tau}}(t,x)),$ (B.3)
for any $(t,x)\in[0,T]\times\mathbb{R}^{d}$. We want to show that for all $\tau\geq 0$
$\left\|\alpha^{\tau}-\alpha^{\tau,\diamond}\right\|_{{\mu^{\tau},\alpha^{\tau}}}\geq c\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\|_{{\mu^{\tau},\alpha^{\tau}}},$ (B.4)
for some constant $c>0$. In order to show (B.4), we assume to the contrary that there exists an increasing sequence $\tau_{k}\to\infty$ such that
$\left\|\alpha^{\tau_{k}}-\alpha^{\tau_{k},\diamond}\right\|_{\mu^{\tau_{k}},\alpha^{\tau_{k}}}\leq\frac{1}{k}\|\alpha^{\tau_{k}}-\alpha^{\mu^{\tau_{k}},*}\|_{\mu^{\tau_{k}},\alpha^{\tau_{k}}}.$ (B.5)
With the short notation
$\alpha_{k}:=\alpha^{\tau_{k}},\,\mu^{k}:=\mu^{\tau_{k}},\,\alpha_{k}^{*}=\alpha^{\tau_{k},*},\,V_{k}:=V^{\mu^{\tau_{k}},\alpha^{\tau_{k}}},\,V_{k}^{*}:=V^{\mu^{k},\alpha_{k}^{*}},\,\alpha_{k}^{\diamond}:=\alpha^{\tau_{k},\diamond},\,\left\|\cdot\right\|_{k}:=\left\|\cdot\right\|_{\mu^{\tau_{k}},\alpha^{\tau_{k}}},$ (B.6)
the inequality above becomes $\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}\leq\frac{1}{k}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}$. With this condition, we can show $\forall\,t\in[0,T]$
$\limsup_{k\to\infty}\int_{\mathbb{R}^{d}}(V_{k}(t,x)-V_{k}^{*}(t,x))\rho_{0}(x)\,\mathrm{d}x=0.$ (B.7)
Note that $V_{k}-V_{k}^{*}$ is always non-negative. We leave the proof for (B.7) to Lemma B.1 in Appendix B.2. This proof is motivated by [55, Theorem 6.1]. After showing (B.7), we let $t=0$ and obtain $\lim_{k\to\infty}(J^{\mu^{k}}[\alpha_{k}]-J^{\mu^{k}}[\alpha_{k}^{*}])=0$. By Lemma A.10, we get
@@
Next, we want to show
@@
By definition of local optimal (B.3), for any $(t,x)\in[0,T]\times\mathbb{R}^{d}$, $\alpha_{k}^{\diamond}(t,x)$ maximizes the map
$\alpha\mapsto H(t,x,\mu^{k}_{t},\alpha,-\nabla_{x}V_{k}(t,x),-\nabla_{x}^{2}V_{k}(t,x)).$
Since $\sigma$ does not depend on $\alpha$, for any fixed $(t,x)$, we view $\alpha_{k}^{\diamond}(t,x)$ as an implicit function of $p=-\nabla_{x}V_{k}(t,x)$, given by the critical point equation
@@
By the optimality of $\alpha_{k}^{*}$ (under distribution $\mu^{k}$), for any $(t,x)\in[0,T]\times\mathbb{R}^{d}$, $\alpha_{k}^{*}(t,x)$ maximizes the map
$\alpha\mapsto H(t,x,\mu^{k}_{t},\alpha,-\nabla_{x}V_{k}^{*}(t,x),-\nabla_{x}^{2}V_{k}^{*}(t,x)).$
Therefore, the implicit function evaluated at $p=-\nabla_{x}V_{k}^{*}(t,x)$ gives $\alpha_{k}^{*}(t,x)$. Next, we want to show that this implicit function is Lipschitz in $p$ with Lipschitz constant $K/\lambda_{H}$. We temporally denote this implicit function by $\alpha(p)$. Computing the Jacobian of the critical point equation $\nabla_{\alpha}H(t,x,\mu^{k}_{t},\alpha(p),p)=0$ with respect to $p\in\mathbb{R}^{d}$, we obtain
$\nabla_{\alpha}^{2}H(t,x,\mu^{k}_{t},\alpha(p),p)\,\nabla_{p}\alpha(p)+\nabla_{\alpha}b(t,x,\mu^{k}_{t},\alpha(p))^{\top}=0,$
which implies
$\nabla_{p}\alpha(p)=-\left(\nabla_{\alpha}^{2}H(t,x,\mu^{k}_{t},\alpha(p),p)\right)^{-1}\cdot\nabla_{\alpha}b(t,x,\mu^{k}_{t},\alpha(p))^{\top}=0.$
Since $H$ is $\lambda_{H}$-strongly concave in $\alpha$, $\left(\nabla_{\alpha}^{2}H(t,x,\mu^{k}_{t},\alpha(p),p)\right)^{-1}$ is positive definite with operator norm less than $\frac{1}{\lambda_{H}}$. Together with $\left|\nabla_{\alpha}b(t,x,\mu^{k}_{t},\alpha(p))\right|\leq K$, we obtain $\left|\nabla_{p}\alpha(p)\right|\leq K/\lambda_{H}$. This further implies
$\left|\alpha_{k}^{\diamond}(t,x)-\alpha_{k}^{*}(t,x)\right|\leq\dfrac{K}{\lambda_{H}}\left|\nabla_{x}V_{k}(t,x)-\nabla_{x}V_{k}^{*}(t,x)\right|.$
Therefore, (B.9) holds.
Next, we want to show
$\left\|\nabla_{x}V_{k}-\nabla_{x}V_{k}^{*}\right\|_{k}\leq C\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}^{1+\chi}$ (B.10)
where $\chi=\frac{2}{d+5}>0$. We leave the proof for (B.10) to Lemma B.2 in Appendix B.3. The result is motivated by (A.18). With (B.10), we have
$\quad\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}\leq\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}+\left\|\alpha_{k}^{\diamond}-\alpha_{k}^{*}\right\|_{k}\leq\frac{1}{k}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}+\dfrac{K}{\lambda_{H}}\left\|\nabla_{x}V_{k}-\nabla_{x}V_{k}^{*}\right\|_{k}$
$\leq\frac{1}{k}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}+C\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}^{1+\chi},$
where the second inequality is due to (B.5) and (B.9); the third inequality is due to (B.10). This contradicts to (B.8) where the contradiction comes from the assumption (B.5). Therefore, (B.4) holds, and
@@
$\geq c\left\|\alpha^{\tau}-\alpha^{\mu,*}\right\|_{\mu^{\tau},\alpha^{\tau}}^{2}\geq c_{a}\left(J^{\mu^{\tau}}[\alpha^{\tau}]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]\right)=c_{a},\mathcal{L}_{a}^{\tau}$
where the first inequality is because of Assumption 4.2, the second inequality is due to (B.4), and the third inequality comes from Lemma A.11.
We preserve the term
$(a\text{IV})=\frac{1}{2}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}.$
@@
$(a\text{V})$ $=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\nabla_{\alpha}b(t,x,\mu_{t},\alpha^{\tau}(t,x))^{\top}\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x)-\mathcal{G}^{\tau}(t,x)\right)\right|^{2}\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\,\mathrm{d}x\,\mathrm{d}t$ (B.12)
$\leq K^{2}\left\|\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}.$
Combining the estimations for $(a\text{III})$ (B.11), $(a\text{IV})$, and $(a\text{V})$, we obtain
@@
Step 2. Next, we estimate $(a\text{II})$. Since $\alpha^{\mu^{\tau},*}$ minimizes $J^{\mu^{\tau}}[\cdot]$, we have
$(a\text{II})=\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]\right)\Big|_{\alpha=\alpha^{\tau}}=\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\prime}]\right)\Big|_{\alpha=\alpha^{\tau},\alpha^{\prime}=\alpha^{\mu^{\tau},*}}$ (B.14)
Next, we want to show
This inequality demonstrate the effect of the OTGP flow to the actor loss function. The proof for (B.15) is deferred to Lemma B.3 in Appendix B.4. Combining (B.14), (B.15), we get
$(a\text{II})\leq C\beta_{\mu}\big\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\,\big\|_{\mu^{\tau},\alpha^{\tau}}^{2}\leq\beta_{\mu}C_{a}\left(J^{\mu^{\tau}}[\alpha^{\tau}]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]\right)=\beta_{\mu}C_{a}\mathcal{L}_{a}^{\tau},$ (B.16)
where the second inequality is due to Lemma A.10.
Finally, we combine the estimations for $(a\text{I})$ in (B.13) and $(a\text{II})$ in (B.16), we get
$\partial_{\tau}\mathcal{L}^{\tau}_{a}$ $\leq-c_{a}\beta_{a}\mathcal{L}_{a}^{\tau}-\frac{1}{2}\beta_{a}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}$
$\quad+\beta_{a}K^{2}\,\big\|\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\big\|_{\mu^{\tau},\alpha^{\tau}}^{2}+C_{a}\beta_{\mu}\mathcal{L}_{a}^{\tau}.$
@@ B.2 Convergence of the gap for value function
Let the assumptions of Theorem 4.4 hold. Assume that (cf. (B.6))
$\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}\leq\frac{1}{k}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}.$ (B.17)
Then,
$\limsup_{k\to\infty}\int_{\mathbb{R}^{d}}(V_{k}(t,x)-V_{k}^{*}(t,x))\rho_{0}(x)\,\mathrm{d}x=0\quad\forall t\in[0,T].$ (B.18)
@@
The proof this this lemma is long, so we outline the strategy here. We split the proof into 7 steps. In Step 1, we assume (to the contrary that) (B.18) does not hold for some $\bar{t}$ and we will restrict our analysis within a small time interval $t_{-},t_{+}]$ which contains $\bar{t}$. Then we apply the doubling of variable technique and define a function $\Phi_{k}$, with maximum denoted by $(t_{k},x_{k},s_{k},y_{k})$. In Step 2, we give some estimations for $(t_{k},x_{k},s_{k},y_{k})$, where Schauder estimation for the value function is applied. In Step 3, we show that $t_{k}$ and $s_{k}$ are not close to $t_{-}$ or $t_{+}$. In Step 4, we add a perturbation to $\Phi_{k}$ and obtain $\widehat{\Phi}_{k}$, whose maximum is denoted by $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$. In Step 5, we obtain a critical point system for this maximum. In Step 6, we conduct some estimation for the critical point system. In Step 7, we take integration among all the local perturbation in Step 4 and obtain a contradiction.
Figure 9: Directed graph describing the dependence of parameters.
Throughout the proof, we will frequently use the properties of Moreau envelope, which are presented in Section A.6. Figure 9 is a directed graph that describes the dependence of parameters in the proof. For example, an arrow from $r_{3}$ to $\iota$ indicates that the choice of $\iota$ depends on $r_{3}$ and potentially also on $r_{3}$’s root nodes including $\mu$, $\varepsilon$, $\lambda$, $R_{2}$ etc. Throughout the proof, we use $C$, $c$ to denote a positive constant that only depends on $d$, $K$, $T$, $\sigma_{0}$, $\lambda_{H}$, and is independent of $k$ and all the parameters in Figure 9. We may record some of these constants as $C_{1}$, $c_{1}$ etc. in order to decide other parameters.
Step 1. We reformulate the problem in this step. We define
@@
Then, by the optimality of $V_{k}^{*}$, $h_{k}(t)\geq 0$ for all $t\in[0,T]$ and $k\geq 1$. The quadratic growth and local Lipschitz property for the value function in Lemma A.12 implies that $\{h_{k}(t)\}_{k=1}^{\infty}$ is uniformly bounded and uniformly Lipschitz. By the Arzelà–Ascoli theorem, $\{h_{k}(t)\}_{k=1}^{\infty}$ has a subsequence that converges uniformly. Therefore, it suffices to show any uniform convergent subsequence of $\{h_{k}(t)\}_{k=1}^{\infty}$ must converge to 0. We assume to the contrary that $\{h_{k}(t)\}_{k=1}^{\infty}$ has a subsequence (without changing the index $k$) that converges uniformly to a nonzero function $h(t)$. Then $h(t)\geq 0$, $h(T)=0$, and $h(t)$ is Lipschitz continuous. We remark that the factor $1/k$ in assumption (B.17) is not essential and can be replaced by any positive sequence that decreases to 0. Thus, retaining the same index notation for $k$ causes no loss of generality.
Since $h(t)$ is Lipschitz, nonzero, and $h(T)=0$, $t_{+}:=\inf\{t\in[0,T]~|~h(s)=0~~~\forall s\in[t,T]\}$ is well defined and $t_{+}>0$. Then, we fix a constant $\Delta t\in(0,t_{+}\wedge 1)$ and denote $t_{-}:=t_{+}-\Delta t$, $I_{t}:=[t_{-},t_{+}]$. The value of $\Delta t$ will be specified later. Next, we pick $\bar{t}\in(t_{-},t_{+})$ s.t. $h(\bar{t})>0$ and denote
@@
Then, we subtract another subsequence of $\{h_{k}(t)\}_{k=1}^{\infty}$ (without changing the index $k$) s.t.
$\int_{\mathbb{R}^{d}}\left(V_{k}(\bar{t},x)-V_{k}^{*}(\bar{t},x)\right)\rho_{0}(x)\,\mathrm{d}x\geq\frac{8}{3}\eta\quad\forall k\geq 1.$
Since $|V_{k}|$ and $|V_{k}^{*}|$ grow at most quadratically in $|x|$ (see Lemma A.12) and $\rho_{0}(x)$ decays exponentially in $|x|$, we can find $R_{1}>0$ and $|\bar{x}_{k}|\leq R_{1}$ for each $k\geq 1$ s.t.
$V_{k}(\bar{t},\bar{x}_{k})-V_{k}^{*}(\bar{t},\bar{x}_{k})\geq\frac{7}{3}\eta\quad\forall k\geq 1.$ (B.19)
Otherwise, we get a contradiction when $R_{1}$ is large:

### v2

@@ Appendix B Proofs for the actor
The proof for Proposition 3.1 is the same as [61, Proposition 1], where we show
@@
Proof of Theorem 4.4.
We decompose the derivative (in $\tau$) of $\mathcal{L}_{a}^{\tau}=J^{\mu^{\tau}}[\alpha^{\tau}]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]$ into two parts
@@
addressing the dependence on $\alpha^{\tau}$ and $\mu^{\tau}$ separately.
@@
$=-\beta_{a}\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla_{\alpha}H\left(t,x,\mu^{\tau}_{t},\alpha^{\tau}(t,x),-\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x)\right)^{\top}$
$\hskip 72.26999pt\nabla_{\alpha}H\left(t,x,\mu^{\tau}_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x)\right)\,\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\,\mathrm{d}x\,\mathrm{d}t$
@@
Firstly, we show $(a\text{III})\geq c\mathcal{L}_{a}^{\tau}$. We start with a technical definition. For any $\tau\geq 0$, $(t,x)\in[0,T]\times\mathbb{R}^{d}$, define the local optimal control $\alpha^{\tau,\diamond}$ as
$\alpha^{\tau,\diamond}(t,x):=\operatorname*{arg\,max}_{a\in\mathbb{R}^{n}}H(t,x,\mu_{t}^{\tau},a,-\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x),-\nabla_{x}^{2}V^{\mu^{\tau},\alpha^{\tau}}(t,x)).$ (B.3)
We want to show that, there exists some constant $c>0$ such that
$\left\|\alpha^{\tau}-\alpha^{\tau,\diamond}\right\|_{{\mu^{\tau},\alpha^{\tau}}}\geq c\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\|_{{\mu^{\tau},\alpha^{\tau}}},\ \forall\tau\geq 0.$ (B.4)
We prove (B.4) by contradiction, assuming that there exists an increasing sequence $\tau_{k}\to\infty$ such that
$\left\|\alpha^{\tau_{k}}-\alpha^{\tau_{k},\diamond}\right\|_{\mu^{\tau_{k}},\alpha^{\tau_{k}}}\leq\frac{1}{k}\|\alpha^{\tau_{k}}-\alpha^{\mu^{\tau_{k}},*}\|_{\mu^{\tau_{k}},\alpha^{\tau_{k}}},\ \forall k\in\mathbb{N}.$ (B.5)
With shorthand notations
$\alpha_{k}:=\alpha^{\tau_{k}},\,\mu^{k}:=\mu^{\tau_{k}},\,\alpha_{k}^{*}=\alpha^{\mu^{k},*},\,V_{k}:=V^{\mu^{\tau_{k}},\alpha^{\tau_{k}}},\,V_{k}^{*}:=V^{\mu^{k},\alpha_{k}^{*}},\,\alpha_{k}^{\diamond}:=\alpha^{\tau_{k},\diamond},\,\left\|\cdot\right\|_{k}:=\left\|\cdot\right\|_{\mu^{\tau_{k}},\alpha^{\tau_{k}}},$ (B.6)
the inequality above becomes $\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}\leq\frac{1}{k}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}$. With this condition, we can show that
$\limsup_{k\to\infty}\int_{\mathbb{R}^{d}}(V_{k}(t,x)-V_{k}^{*}(t,x))\rho_{0}(x)\,\mathrm{d}x=0,\ \forall t\in[0,T],$ (B.7)
with its proof (motivated by [56, Theorem 6.1]) left to Lemma B.1 in Appendix B.2. Since $V_{k}-V_{k}^{*}$ is non-negative by definition, setting $t=0$ yields $\lim_{k\to\infty}(J^{\mu^{k}}[\alpha_{k}]-J^{\mu^{k}}[\alpha_{k}^{*}])=0$. By Lemma A.10, we get
@@
As an intermediate step toward reaching contradiction, we prove that
@@
Since $\sigma$ does not depend on $\alpha$, by definition  (B.3), we view $\alpha_{k}^{\diamond}(t,x)$ as an implicit function of $p=-\nabla_{x}V_{k}(t,x)$ for any fixed tuple of $(t,x)$, with the functional relationship determined by the critical point equation
@@
By the optimality of $\alpha_{k}^{*}$ (under measure $\mu^{k}$), for any $(t,x)\in[0,T]\times\mathbb{R}^{d}$, $\alpha_{k}^{*}(t,x)$ maximizes the mapping $\alpha\mapsto H(t,x,\mu^{k}_{t},\alpha,-\nabla_{x}V_{k}^{*}(t,x),-\nabla_{x}^{2}V_{k}^{*}(t,x)).$ Therefore, the same implicit function evaluated at $p=-\nabla_{x}V_{k}^{*}(t,x)$ provides $\alpha_{k}^{*}(t,x)$. Naturally, showing (B.9) reduces to showing that this implicit function is Lipschitz in $p$ with Lipschitz constant $K/\lambda_{H}$. Recall that $H$ is $\lambda_{H}$-strongly concave in $\alpha$. By the implicit function theorem, the continuously differentiable implicit function $\alpha(p)$ globally exists. Computing its Jacobian with respect to $p\in\mathbb{R}^{d}$ yields
$\nabla_{p}\alpha(p)=-\left(\nabla_{\alpha}^{2}H(t,x,\mu^{k}_{t},\alpha(p),p)\right)^{-1}\cdot\nabla_{\alpha}b(t,x,\mu^{k}_{t},\alpha(p)).$
Since $\left|\nabla_{\alpha}b\right|\leq K$ and $\|\left(\nabla_{\alpha}^{2}H(t,x,\mu^{k}_{t},\alpha(p),p)\right)^{-1}\|_{2}\leq\frac{1}{\lambda_{H}}$, we obtain $\left|\nabla_{p}\alpha(p)\right|\leq K/\lambda_{H}$, which concludes the proof of (B.9).
The final step toward reaching contradiction is to show the superlinear growth
$\left\|\nabla_{x}V_{k}-\nabla_{x}V_{k}^{*}\right\|_{k}\leq C\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}^{1+\chi},$ (B.10)
where $\chi=\frac{2}{d+5}>0$. We leave the proof of (B.10) (motivated by (A.18)) to Lemma B.2 in Appendix B.3.
At this point, we present the contradiction, which proves (B.4). Using (B.5), (B.9) and (B.10),
$\quad\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}\leq\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}+\left\|\alpha_{k}^{\diamond}-\alpha_{k}^{*}\right\|_{k}\leq\frac{1}{k}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}+C\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}^{1+\chi},$
which contradicts (B.8) as $k\to\infty$.
Now, we return to showing $(a\text{III})\geq c\mathcal{L}_{a}^{\tau}$ with the help of (B.4):
@@
$\geq c\left\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\right\|_{\mu^{\tau},\alpha^{\tau}}^{2}\geq c_{a}\left(J^{\mu^{\tau}}[\alpha^{\tau}]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]\right)=c_{a}\mathcal{L}_{a}^{\tau},$
where the first inequality is due to the mean value theorem and Assumption 4.2, and the last inequality comes from Lemma A.11.
@@
$(a\text{V})$ $=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\nabla_{\alpha}b(t,x,\mu_{t},\alpha^{\tau}(t,x))^{\top}\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x)-\mathcal{G}^{\tau}(t,x)\right)\right|^{2}\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\,\mathrm{d}x\,\mathrm{d}t$ (B.12)
$\leq\frac{1}{2}K^{2}\left\|\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}.$
Combining estimations for $(a\text{III})$ (cf. (B.11)), $(a\text{IV})$ and $(a\text{V})$ yields
@@
Step 2. Next, we estimate $(a\text{II})$. Since $\alpha^{\mu^{\tau},*}$ minimizes $J^{\mu^{\tau}}[\cdot]$, by chain rule,
$(a\text{II})=\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]\right)\Big|_{\alpha=\alpha^{\tau}}=\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\prime}]\right)\Big|_{\alpha=\alpha^{\tau},\alpha^{\prime}=\alpha^{\mu^{\tau},*}}.$ (B.14)
We claim that
$\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(J^{\mu^{\tau}}[\alpha]-J^{\mu^{\tau}}[\alpha^{\prime}]\right)\Big|_{\alpha=\alpha^{\tau},\alpha^{\prime}=\alpha^{\mu^{\tau},*}}\leq C\beta_{\mu}\big\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\big\|_{\mu^{\tau},\alpha^{\tau}}^{2},$ (B.15)
the proof of which is deferred to Lemma B.3 in Appendix B.4. This inequality demonstrates the impact of the distribution flow on the actor loss function. Combining (B.14), (B.15) and Lemma A.10 yields
$(a\text{II})\leq C\beta_{\mu}\big\|\alpha^{\tau}-\alpha^{\mu^{\tau},*}\,\big\|_{\mu^{\tau},\alpha^{\tau}}^{2}\leq\beta_{\mu}C_{a}\left(J^{\mu^{\tau}}[\alpha^{\tau}]-J^{\mu^{\tau}}[\alpha^{\mu^{\tau},*}]\right)=\beta_{\mu}C_{a}\mathcal{L}_{a}^{\tau}.$ (B.16)
Finally, combining estimations for $(a\text{I})$ (cf. (B.13)) and $(a\text{II})$ (cf. (B.16)) concludes the proof. ∎
@@ B.2 Convergence of the gap for value function
Under the notations of (B.6), if the conditions of Theorem 4.4 hold and
$\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}\leq\frac{1}{k}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k},\ \forall k\geq 1,$ (B.17)
then we claim that
$\limsup_{k\to\infty}\int_{\mathbb{R}^{d}}(V_{k}(t,x)-V_{k}^{*}(t,x))\rho_{0}(x)\,\mathrm{d}x=0,\ \forall t\in[0,T].$ (B.18)
@@
We prove this lemma by contradiction. Since the proof is very long, we split it into 7 steps, introducing the strategy before diving into details.
In Step 1, we assume the existence of some $\bar{t}$, for which (B.18) fails. We restrict our analysis within a small time interval $t_{-},t_{+}]$ containing $\bar{t}$, apply the doubling of variable technique, and define a function $\Phi_{k}$, with its maximum attained by the tuple $(t_{k},x_{k},s_{k},y_{k})$. In Step 2, we estimate $(t_{k},x_{k},s_{k},y_{k})$ using the Schauder estimation for value functions. In Step 3, we show that both $t_{k}$ and $s_{k}$ are not close to $t_{-}$ or $t_{+}$.
In Step 4, we consider $\widehat{\Phi}_{k}$, a perturbed version of $\Phi_{k}$, whose maximum is attained by $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$. In Step 5, we obtain a critical point system for $\widehat{\Phi}_{k}$ at its maximum. In Step 6, we carry out estimations for the critical point system. In Step 7, we integrate with respect to all local perturbations in Step 4 and reach a contradiction.
Figure 9: A directed graph describing parameter dependencies.
Throughout the proof, we will frequently use the properties of Moreau envelopes presented in Section A.6. Figure 9 describes parameter dependencies in the proof through a directed graph. For example, an arrow from $r_{3}$ to $\iota$ indicates that the choice of $\iota$ potentially depends on $r_{3}$ and its ancestor nodes, including $\mu$, $\lambda$, $\varepsilon$, $R_{2}$, etc. Without specification, $C$ and $c$ denote positive constants that only depend on $d$, $K$, $T$, $\sigma_{0}$, $\lambda_{H}$, which are uniform with respect to $k$ and all the parameters in Figure 9. We may denote some of these constants by $C_{1}$, $c_{1}$, etc., for the specification of other parameters.
Step 1. Firstly, we reformulate the problem. Define
@@
By the optimality of $V_{k}^{*}$, $h_{k}(t)\geq 0,\ \forall t\in[0,T],\ k\geq 1$. By the quadratic growth and local Lipschitz property of value functions in Lemma A.12, $\{h_{k}(t)\}_{k=1}^{\infty}$ is uniformly bounded and uniformly Lipschitz. By the Arzelà–Ascoli theorem, $\{h_{k}(t)\}_{k=1}^{\infty}$ has a subsequence that converges uniformly on $[0,T]$. Therefore, it suffices to show that any uniformly convergent subsequence of $\{h_{k}(t)\}_{k=1}^{\infty}$ must converge to 0.
We prove by contradiction, assuming that $\{h_{k}(t)\}_{k=1}^{\infty}$ has a subsequence that converges uniformly to a nonzero function $h(t)$. For simplicity of notations, we still use $\{h_{k}(t)\}_{k=1}^{\infty}$ to denote this subsequence in the following context without specification. We remark that, the factor $1/k$ in the condition (B.17) can be replaced by any positive sequence that decreases to 0. Hence, using the same index $k$ for the subsequence causes no loss of generality.
Clearly, $h(t)\geq 0$, $h(T)=0$, and $h$ is Lipschitz continuous. Since $h$ is not constantly zero, $t_{+}:=\inf\{t\in[0,T]~|~h(s)=0,\ \forall s\in[t,T]\}$ is well defined and $t_{+}>0$. For a fixed value of $\Delta t\in(0,t_{+}\wedge 1)$, which will be later specified, define $t_{-}:=t_{+}-\Delta t$, $I_{t}:=[t_{-},t_{+}]$. We pick $\bar{t}\in(t_{-},t_{+})$ such that $h(\bar{t})>0$ and define
@@
There exists a further subsequence of the convergent subsequence $\{h_{k}(t)\}_{k=1}^{\infty}$ (which is still denoted by $\{h_{k}(t)\}_{k=1}^{\infty}$), such that
$\int_{\mathbb{R}^{d}}\left(V_{k}(\bar{t},x)-V_{k}^{*}(\bar{t},x)\right)\rho_{0}(x)\,\mathrm{d}x\geq\frac{8}{3}\eta,\ \forall k\geq 1.$
Since $|V_{k}|$ and $|V_{k}^{*}|$ grow at most quadratically in $|x|$ (see Lemma A.12) and $\rho_{0}(x)$ decays exponentially in $|x|$, we claim that: there exists $R_{1}>0$ and $|\bar{x}_{k}|\leq R_{1},\ \forall k\geq 1$ such that
$V_{k}(\bar{t},\bar{x}_{k})-V_{k}^{*}(\bar{t},\bar{x}_{k})\geq\frac{7}{3}\eta,\ \forall k\geq 1.$ (B.19)
Otherwise, if such $R_{1}$ and $\bar{x}_{k}$ do not exist, then for some $k$, $V_{k}(\bar{t},\bar{x}_{k})-V_{k}^{*}(\bar{t},\bar{x}_{k})<\frac{7}{3}\eta$. This implies that